A balloon is filled to a volume of 6.90L at a temperature of 27.1ºC. If the pressure in the balloon is measured to be 2.20 atm, how many moles of gas are contained inside the balloon?
1
Expert's answer
2019-12-05T08:19:47-0500

According to the Clapeyron-Mendeleev equation, the amount of the ideal gas in moles can be expressed as:


"n=\frac{PV}{RT}=\frac{2.20*101300Pa*6.90*10^{-3}m^3}{8.31\frac{J}{mol*K}*300.1K}=0.617mol."


Thus, there were 0.617 moles of this gas.